# Polarimeter to Unify the Corona and Heliosphere (PUNCH)

### v1

@@ 6 Data products and Ground Pipeline
PUNCH data products are generated in the Science Operations Center (SOC), which is described in detail by Hughes et al. (2025) and summarized here. Data are produced as FITS files (Wells, Greisen, and Harten, 1981) in four principal processing Levels, numbered from 0 through 3. Background subtraction is the driving challenge for PUNCH data reduction, and drives precision of all levels of data reduction.
@@ 6.2 Level 1 data
Level 1 pixels have floating-point values that are intended to capture photometric values on the sky, but are still individual exposures from particular Observatories through particular polarizers. The Level 0 to Level 1 steps include: improved square-root decoding (DeForest et al., 2022); destreaking to account for quasi-frame-transfer streaks in camera operation; quartic-polynomial flat-field correction including both detector effects and optical vignetting; cosmic ray and saturated-pixel detection and removal; pointing and distortion-function correction using the starfield; and point-spread function (PSF) regularization (Hughes et al., 2025). The pixel values are calibrated in $B_{\odot}$ units (using the mean photospheric radiance of $7.58\times 10^{6}$ W m-2 SR-1, adjusted to the passband of PUNCH).
@@ 6.3 Level 2 data
The polarized data products are typically 3-D “cubes” of size $w\times h\times 3$, with $w$ being the image width, $h$ being the image height, and the third dimension running across MZP666for Minus, Zero, Plus [60] polarizer angle relative to a reference angle virtual-polarizer triplet images (Appendix A.2), while the unpolarized/clear data products are typically 2-D $w\times h$ images. The MZP virtual polarizer angles are measured relative to the system in which solar north is up; the system is described by DeForest, Seaton, and West (2022) and developed in more detail by Patel, Seaton, and DeForest (2025).
@@ 6.4 Level 3 data
Figure 9 illustrates the fundamental challenge of background subtraction for all heliospheric imagers including PUNCH: the F corona (zodiacal light) and starfield dominate over the signal by 2-3 orders of magnitude in the far field. PUNCH uses the same six-step process for background subtraction as was developed for STEREO/HI (DeForest, Howard, and Tappin, 2011; DeForest et al., 2016), with the advantage of additional lessons learned from that mission, and also the challenge of using spliced mosaic images rather than single exposures from one optical system. The process is summarized here and detailed by Hughes et al. (2025).
@@
The second background is the F corona (zodiacal light), which is quasi-stationary in the heliocentric reference frame and is removed on the basis of a running-minimum fit, corrected for variable noise level effects, in the Level 2 data. Each pixel location in the mosaic frame is fit with a running-minimum polynomial background estimator; these polynomials are conditioned to produce daily F corona models, which are interpolated to provide an F background for each data product, and subtracted in the L3 pipeline. This method makes use of the observed approximate symmetry of the heliospheric dust cloud; it is expected that any dynamic features such as a dust enhancement near Venus (e.g., Jones, Bewsher, and Brown, 2017; Stenborg et al., 2021a) or local dust depletions (Stenborg et al., 2021b) will be recognizable as their dynamics follow Keplerian orbital time scales of months, rather than solar wind time scales of days, and may be removed on that basis.
@@
Following DeForest, Howard, and Tappin (2011), a final step of Fourier filtering is applied to clean the L3 data via motion filtering. The motion-filtering step for PUNCH is sufficiently broad to pass SIRs and other caustic structures that may move $3\times$$5\times$ slower than solar wind itself.
@@ 7.1 Stellar polarimetric observations
The bulk of the stellar linear polarization signal is not due to intrinsic polarization of emitted starlight, but to aspherical grains interacting with the light as it propagates: either in interstellar space (Versteeg et al., 2023) or (less commonly) in local dust clouds (Poeckert, Bastien, and Landstreet, 1979). The polarization of starlight by interstellar dust is interesting in part because it is a measure of the direction and strength of the interstellar magnetic field (Versteeg et al., 2023), and of turbulence in the interstellar medium (Angarita et al., 2024).
@@ 7.2 Asteroid and comet observations
PUNCH does what the other asteroid surveys can’t do - it observes at very low solar elongation and it does so all day every day. In fact, some of the objects predicted to populate this region will always be in the field of view of PUNCH. The ability for PUNCH to discover, track and characterize asteroids and comets depends on the trade between its relatively small imaging aperture, its large field of view and its rapid and constant imaging cadence. There is precedent for observing asteroids and comets with SOHO and STEREO (Knight et al., 2010; Li and Jewitt, 2013), and the larger field of view, smaller pixels and lack of on-board image stacking may permit PUNCH to be a more effective tool for characterization and for discovering new objects.
@@ 7.3 Solar system dust observations
Finally, with its increased sensitivity, PUNCH will likely be able to detect subtle brightness increases due to, for example, asteroids or comets entering the inner heliosphere as well the impact of interstellar dust increasing the brightness in a range of ecliptic longitudes centered at $259^{\circ}$, due to the solar system traveling through the galaxy. The range of longitudes could be as high as $180^{\circ}$ (full-width), but the increase would be highest at the central longitude. Additional searches include finding the dust depletion zone, which begins at 35 $R_{\odot}$ and peaks at 25 $R_{\odot}$ (Stenborg et al., 2021b), or even if there are additional sublimation zones beyond 35 $R_{\odot}$; and characterizing and understanding dust depletions behind CMEs as have been seen with WISPR (Stenborg et al., 2023), or associated with other plasma structures such as SIRs.
@@ Acknowledgements.
PUNCH is the result of a very large scientific and technical community working together over a decade, combining outstanding talent across many fields and heroic effort through many types of adversity. The literally hundreds of individuals, whose dedication and willing support brought PUNCH to fruition, span four continents and dozens of institutions. PUNCH also builds on the major achievements of prior missions, including PSP/WISPR, STEREO/SECCHI, SMEI, SOHO/LASCO, and the Helios three-channel photometer; and would not have happened without participation, support, and expertise from the leadership of those projects. PUNCH owes much of its overall design and scientific concept to early efforts by T.A. Howard, whose early collaborative work was instrumental in co-developing and exploiting methods for deep background-subtraction of heliospheric images; in predicting and promoting the possibility of wide-field 3-D analysis via polarization; and in helping demonstrate the feasibility of heliospheric imaging from below the Van Allen belts. We particularly acknowledge the vendor Tele Vue Optics (Al Nagler) for designing and supplying the optics for the Wide Field Imagers, working well outside Tele Vue’s “comfort zone” of eyepiece design and stepping up to produce a very compact, achromatic, space-qualified system with very low distortion, several important stray-light reducing features, and excellent focus over a very wide field of view. J.B., S.F., S.E.G., and A.M. acknowledge support of the National Center for Atmospheric Research, a major facility sponsored by the National Science Foundation under Cooperative Agreement No. 1852977. D.J.M acknowledges support from the IMAP mission as a part of NASA’s STP mission line (80GSFC19C0027) PUNCH is a heliophysics mission to study the corona, solar wind, and space weather as an integrated system, and is part of NASA’s Explorers program (Contract 80GSFC14C0014).
@@ A.1 Geometric definitions
Direct 2-D image analyses of the corona and heliosphere have typically treated features as if they’re projected onto the tangent plane: because $r_{obs}$ is known and $\varepsilon$ is measured by direct location in the image plane, the related distance $r_{pos}=r_{obs}\tan\varepsilon$ is simple to calculate. Alternatively, in wide field studies (e.g. with STEREO/SECCHI, Howard et al., 2008), features are sometimes considered to be on or near the Thomson surface (e.g. Vourlidas and Howard, 2006; DeForest, Howard, and Tappin, 2013), placing them at the line-of-sight impact parameter distance $d=r_{obs}\sin\varepsilon$ from the Sun.
@@ A.2 Photometric and polarimetric definitions
Polarimetry may be (and has been) described in multiple ways by multiple authors; the fundamentals of the Stokes parameter formulation are covered in standard textbooks (e.g. Hecht and Zajac, 1974; Born and Wolf, 1999). The polarization signal expected from the far-field K corona, and the PUNCH approach to measuring it, are developed in detail by DeForest, Howard, and Tappin (2013) and DeForest, Seaton, and West (2022), respectively, and summarized here.
@@
Thomson scattering polarizes light perpendicular to the scattering plane (Figure 13), by projection of the electric field from the incident ray into the plane perpendicular to the scattered ray. This means that the radiance $B_{R}$ viewed through a radially-aligned linear polarizer is less than the radiance $B_{T}$ viewed through a tangentially-aligned polarizer (i.e. one oriented $90^{\circ}$ to the solar-radial direction). The quantity $B_{T}-B_{R}$ (“tangential polarized brightness”) is similar to the Stokes parameter $Q$, with the notable difference that $Q$ is defined in the fixed reference frame of a laboratory, while the orientation of the polarizers for $B_{T}$ and $B_{R}$ varies around the image plane. We refer to $B_{T}-B_{R}$ as $pB$ and identify it with the $^{\perp}pB$ defined by DeForest, Seaton, and West (2022). The appendix to that article, incidentally, includes a discussion of the (confusing) history of $pB$ nomenclature, including the contrast between $^{\perp}pB$ and the related but distinct quantity $^{\circ}pB~\equiv~\sqrt{Q^{2}+U^{2}}$ (“total linear polarized brightness”). Both quantitites have been called $pB$ in the historical literature by different (or sometimes even the same) authors; to avoid confusion, we encourage authors to define $pB$ explicitly in every article that uses the term. The quantity $^{\circ}pB$ has also been called $L$ (Rachmeler et al., 2013), a usage we encourage for its distinctness from $pB$.
By analogy to the Stokes parameters, there exists a parameter $pB^{\prime}$ that is conjugate to $pB$ and describes linear polarization diagonal to the radial direction on the image plane. $pB^{\prime}$ is both predicted and observed to be zero in coronal Thomson-scattered light (e.g., Filippov, Molodensky, and Koutchmy, 1994). It is omitted from PUNCH L3 published data products, although the data reduction pipeline can produce $pB^{\prime}$ image products as a check on data quality.
PUNCH does not measure $B_{R}$ and $B_{T}$ directly. Instead, each instrument measures polarization triplets: sequences of three exposures taken through linear polarizers oriented $60^{\circ}$ apart. The theory has been known for some time (Fesenkov, 1935; Öhman, 1947) and is recapitulated, with an eye toward modern application, by DeForest, Seaton, and West (2022). We refer to the polarization-triplet system as MZP, and to each brightness value in a triplet as $B_{M}$, $B_{Z}$, or $B_{P}$ respectively. PUNCH collects MZP triplets in each instrument’s reference frame on-orbit, then mathematically transforms them to MZP in rectified solar coordinates. This requires accounting not only for spacecraft rotation angle, but also esoterica such as projection effects that change the effective angle of each polarizer across the image plane (Patel, Seaton, and DeForest, 2025).
@@ Appendix B Polarization
PUNCH polarimetry is specifically designed to locate objects in three dimensions, via the physics of Thomson scattering. The theory has been developed to varying degrees of detail by many authors, from Minnaert (1930) and van de Hulst (1950), through Billings (1966), to modern works including Mierla et al. (2009) and Howard and Tappin (2009). It has been demonstrated by several authors including DeForest, de Koning, and Elliott (2017) and Mierla et al. (2022). PUNCH observes features that are quite far from the Sun and therefore polarimetric analysis may be simplified considerably via the point-sun approximation (DeForest, Howard, and Tappin, 2013), which is important given the complexity of large distributed objects such as CMEs (Howard et al., 2013). In this Appendix we very briefly discuss this elementary theory of polarimetry to locate features, and refer to both prior articles that lay a more comprehensive foundation and other articles in this collection, which explore the details of distributed object polarimetry in the context of PUNCH.
@@
where $p\equiv pB/tB$ is the degree of linear polarization from Thomson scattering. Both the $\arccos$ and the square root have sign ambiguities, but the positive branch of the square root is forced by the geometry of the scattering plane, leaving only the sign ambiguity of the $\arccos$ function. That ambiguity leads to a split between a “real” and a “ghost” location along the line of sight, for a feature seen in any one observation. The two locations may be distinguished by their behavior over time in multiple image sets: small ejecta from the Sun are known to follow radial trajectories, and the ghost location follows a clearly nonphysical trajectory (DeForest, Howard, and Tappin, 2013; Gibson et al., 2025).
@@
Figure 14 illustrates simple 3D inversion, applicable to compact features that are well isolated and fully background subtracted. Reading off the polarization ratio yields an estimate of the $\tau$ angle (complement of the scattering angle $\chi$). From that and the elongation $\varepsilon$ of the feature, one may infer the out-of-sky-plane angle $\xi$, as shown in Figure 13. This compact-feature case, for features at least several $r_{\odot}$ from the Sun itself, has been developed analytically by several authors including DeForest, Howard, and Tappin (2013). This compact-feature approximation may be used even for portions of large features (whose size is comparable to the distance from the observer or the Sun), because the 2-D leading edge of any large feature must necessarily be spatially compact along the line of sight (Howard et al., 2013). However, all coronal features exist against the background of myriad other K corona (or solar wind) features along each line of sight, and that “K background” must be considered and removed in addition to the backgrounds from stray light, F corona, and stellar backgrounds described in Section 6.4 (e.g., van de Hulst, 1950; DeForest, de Koning, and Elliott, 2017).

### v2

@@ 6 Data products and Ground Pipeline
PUNCH data products are generated in the Science Operations Center (SOC), which is described in detail by Hughes et al. (2025) and summarized here. Data are produced as FITS files (Wells et al., 1981) in four principal processing Levels, numbered from 0 through 3. Background subtraction is the driving challenge for PUNCH data reduction, and drives precision of all levels of data reduction.
@@ 6.2 Level 1 data
Level 1 pixels have floating-point values that are intended to capture photometric values on the sky, but are still individual exposures from particular Observatories through particular polarizers. The Level 0 to Level 1 steps include: improved square-root decoding (DeForest et al., 2022a); destreaking to account for quasi-frame-transfer streaks in camera operation; quartic-polynomial flat-field correction including both detector effects and optical vignetting; cosmic ray and saturated-pixel detection and removal; pointing and distortion-function correction using the starfield; and point-spread function (PSF) regularization (Hughes et al., 2025). The pixel values are calibrated in $B_{\odot}$ units (using the mean photospheric radiance of $7.58\times 10^{6}$ W m-2 SR-1, adjusted to the passband of PUNCH).
@@ 6.3 Level 2 data
The polarized data products are typically 3-D “cubes” of size $w\times h\times 3$, with $w$ being the image width, $h$ being the image height, and the third dimension running across MZP666for Minus, Zero, Plus [60] polarizer angle relative to a reference angle virtual-polarizer triplet images (Appendix A.2), while the unpolarized/clear data products are typically 2-D $w\times h$ images. The MZP virtual polarizer angles are measured relative to the system in which solar north is up; the system is described by DeForest et al. (2022b) and developed in more detail by Patel et al. (2025).
@@ 6.4 Level 3 data
Figure 10 illustrates the fundamental challenge of background subtraction for all heliospheric imagers including PUNCH: the F corona (zodiacal light) and starfield dominate over the signal by 2-3 orders of magnitude in the far field. PUNCH uses the same six-step process for background subtraction as was developed for STEREO/HI (DeForest et al., 2011, 2016), with the advantage of additional lessons learned from that mission, and also the challenge of using spliced mosaic images rather than single exposures from one optical system. The process is summarized here and detailed by Hughes et al. (2025).
@@
The second background is the F corona (zodiacal light), which is quasi-stationary in the heliocentric reference frame and is removed on the basis of a running-minimum fit, corrected for variable noise level effects, in the Level 2 data. Each pixel location in the mosaic frame is fit with a running-minimum polynomial background estimator; these polynomials are conditioned to produce daily F corona models, which are interpolated to provide an F background for each data product, and subtracted in the L3 pipeline. This method makes use of the observed approximate symmetry of the heliospheric dust cloud; it is expected that any dynamic features such as a dust enhancement near Venus (e.g., Jones et al., 2017; Stenborg et al., 2021b) or local dust depletions (Stenborg et al., 2021a) will be recognizable as their dynamics follow Keplerian orbital time scales of months, rather than solar wind time scales of days, and may be removed on that basis.
@@
Following DeForest et al. (2011), a final step of Fourier filtering is applied to clean the L3 data via motion filtering. The motion-filtering step for PUNCH is sufficiently broad to pass SIRs and other caustic structures that may move $3\times$$5\times$ slower than solar wind itself.
@@ 7.1 Stellar polarimetric observations
The bulk of the stellar linear polarization signal is not due to intrinsic polarization of emitted starlight, but to aspherical grains interacting with the light as it propagates: either in interstellar space (Versteeg et al., 2023) or (less commonly) in local dust clouds (Poeckert et al., 1979). The polarization of starlight by interstellar dust is interesting in part because it is a measure of the direction and strength of the interstellar magnetic field (Versteeg et al., 2023), and of turbulence in the interstellar medium (Angarita et al., 2024).
@@ 7.2 Asteroid and comet observations
PUNCH observes continuously in a part of the sky that is historically difficult: areas with low solar elongation. Many of the objects predicted to populate this region are always in the PUNCH FOV. The ability for PUNCH to discover, track and characterize asteroids and comets depends on the trade between its relatively small imaging aperture, its large field of view and its rapid and constant imaging cadence. There is precedent for observing asteroids and comets with SOHO and STEREO (Knight et al., 2010; Li and Jewitt, 2013), and the larger field of view, smaller pixels and lack of on-board image stacking may permit PUNCH to be a more effective tool for characterization and for discovering new objects.
@@ 7.3 Solar system dust observations
Finally, with its increased sensitivity, PUNCH will likely be able to detect subtle brightness increases due to, for example, asteroids or comets entering the inner heliosphere as well the impact of interstellar dust increasing the brightness in a range of ecliptic longitudes centered at $259^{\circ}$, due to the solar system traveling through the galaxy. The range of longitudes could be as high as $180^{\circ}$ (full-width), but the increase would be highest at the central longitude. Additional searches include finding the dust depletion zone, which begins at 35 $R_{\odot}$ and peaks at 25 $R_{\odot}$ (Stenborg et al., 2021a), or even if there are additional sublimation zones beyond 35 $R_{\odot}$; and characterizing and understanding dust depletions behind CMEs as have been seen with WISPR (Stenborg et al., 2023), or associated with other plasma structures such as SIRs.
@@ Acknowledgements.
PUNCH is the result of a very large scientific and technical community working together over a decade, combining outstanding talent across many fields and heroic effort through many types of adversity. The hundreds of individuals, whose dedication and willing support brought PUNCH to fruition, span four continents and dozens of institutions. PUNCH builds on the major achievements of prior missions, including PSP/WISPR, STEREO/SECCHI, Coriolis/SMEI, SOHO/LASCO, and the Helios three-channel photometer. PUNCH owes much of its overall design and scientific concept to early efforts by T.A. Howard, whose work was instrumental in co-developing and exploiting methods for deep background-subtraction of heliospheric images; in predicting and promoting the possibility of wide-field 3-D analysis via polarization; and in helping demonstrate the feasibility of heliospheric imaging from below the Van Allen belts. We particularly acknowledge Tele Vue Optics and remember their founder, Al Nagler (1935-2025), for designing and supplying the optics for the Wide Field Imagers, working well outside Tele Vue’s “comfort zone” of eyepiece design, and stepping up to produce a very compact, achromatic, space-qualified system with very low distortion, several important stray-light reducing features, and excellent focus over a very wide field of view. PUNCH/WFI is but one of Al Nagler’s many contributions to the astronomy, spaceflight, and scientific communities, and he is both sorely missed and fondly remembered. We also gratefully acknowledge C. Eyles for early consultation on heliospheric imager design based on his experience with SMEI; the SwRI Internal Research program for much early support that led to the successful PUNCH proposal; J. Burch for consultation and institutional support throughout the project; and J. Andrews for invaluable advice on project management. J.B., S.F., S.E.G., and A.M. acknowledge support of the National Center for At- mospheric Research, a major facility sponsored by the National Science Foundation under Cooperative Agreement No. 1852977. D.J.M acknowledges support from the IMAP mission as a part of NASA’s STP mission line (80GSFC19C0027). PUNCH is a heliophysics mission to study the corona, solar wind, and space weather as an integrated system, and is part of NASA’s Explorers program (Contract 80GSFC14C0014).
@@ A.1 Geometric definitions
Direct 2-D image analyses of the corona and heliosphere have typically treated features as if they’re projected onto the tangent plane: because $r_{obs}$ is known and $\varepsilon$ is measured by direct location in the image plane, the related distance $r_{pos}=r_{obs}\tan\varepsilon$ is simple to calculate. Alternatively, in wide field studies (e.g. with STEREO/SECCHI, Howard et al., 2008), features are sometimes considered to be on or near the Thomson surface (e.g. Vourlidas and Howard, 2006; DeForest et al., 2013b), placing them at the line-of-sight impact parameter distance $d=r_{obs}\sin\varepsilon$ from the Sun.
@@ A.2 Photometric and polarimetric definitions
Polarimetry may be (and has been) described in multiple ways by multiple authors; the fundamentals of the Stokes parameter formulation are covered in standard textbooks (e.g. Hecht and Zajac, 1974; Born and Wolf, 1999). The polarization signal expected from the far-field K corona, and the PUNCH approach to measuring it, are developed in detail by DeForest et al. (2013b) and DeForest et al. (2022b), respectively, and summarized here.
@@
Thomson scattering polarizes light perpendicular to the scattering plane (Figure 18), by projection of the electric field from the incident ray into the plane perpendicular to the scattered ray. This means that the radiance $B_{R}$ viewed through a radially-aligned linear polarizer is less than the radiance $B_{T}$ viewed through a tangentially-aligned polarizer (i.e. one oriented $90^{\circ}$ to the solar-radial direction). The quantity $B_{T}-B_{R}$ (“tangential polarized brightness”) is similar to the Stokes parameter $Q$, with the notable difference that $Q$ is defined in the fixed reference frame of a laboratory, while the orientation of the polarizers for $B_{T}$ and $B_{R}$ varies around the image plane. We refer to $B_{T}-B_{R}$ as $pB$ and identify it with the $^{\perp}pB$ defined by DeForest et al. (2022b). The appendix to that article, incidentally, includes a discussion of the (confusing) history of $pB$ nomenclature, including the contrast between $^{\perp}pB$ and the related but distinct quantity $^{\circ}pB~\equiv~\sqrt{Q^{2}+U^{2}}$ (“total linear polarized brightness”). Both quantitites have been called $pB$ in the historical literature by different (or sometimes even the same) authors; to avoid confusion, we encourage authors to define $pB$ explicitly in every article that uses the term. The quantity $^{\circ}pB$ has also been called $L$ (Rachmeler et al., 2013), a usage we encourage for its distinctness from $pB$.
By analogy to the Stokes parameters, there exists a parameter $pB^{\prime}$ that is conjugate to $pB$ and describes linear polarization diagonal to the radial direction on the image plane. $pB^{\prime}$ is both predicted and observed to be zero in coronal Thomson-scattered light (e.g., Filippov et al., 1994). It is omitted from PUNCH L3 published data products, although the data reduction pipeline can produce $pB^{\prime}$ image products as a check on data quality.
PUNCH does not measure $B_{R}$ and $B_{T}$ directly. Instead, each instrument measures polarization triplets: sequences of three exposures taken through linear polarizers oriented $60^{\circ}$ apart. The theory has been known for some time (Fesenkov, 1935; Öhman, 1947) and is recapitulated, with an eye toward modern application, by DeForest et al. (2022b). We refer to the polarization-triplet system as MZP, and to each brightness value in a triplet as $B_{M}$, $B_{Z}$, or $B_{P}$ respectively. PUNCH collects MZP triplets in each instrument’s reference frame on-orbit, then mathematically transforms them to MZP in rectified solar coordinates. This requires accounting not only for spacecraft rotation angle, but also esoterica such as projection effects that change the effective angle of each polarizer across the image plane (Patel et al., 2025).
@@ Appendix B Polarization
PUNCH polarimetry is specifically designed to locate objects in three dimensions, via the physics of Thomson scattering. The theory has been developed to varying degrees of detail by many authors, from Minnaert (1930) and van de Hulst (1950), through Billings (1966), to modern works including Mierla et al. (2009) and Howard and Tappin (2009). It has been demonstrated by several authors including DeForest et al. (2017) and Mierla et al. (2022). PUNCH observes features that are quite far from the Sun and therefore polarimetric analysis may be simplified considerably via the point-sun approximation (DeForest et al., 2013b), which is important given the complexity of large distributed objects such as CMEs (Howard et al., 2013). In this Appendix we very briefly discuss this elementary theory of polarimetry to locate features, and refer to both prior articles that lay a more comprehensive foundation and other articles in this collection, which explore the details of distributed object polarimetry in the context of PUNCH.
@@
where $p\equiv pB/tB$ is the degree of linear polarization from Thomson scattering. Both the $\arccos$ and the square root have sign ambiguities, but the positive branch of the square root is forced by the geometry of the scattering plane, leaving only the sign ambiguity of the $\arccos$ function. That ambiguity leads to a split between a “real” and a “ghost” location along the line of sight, for a feature seen in any one observation. The two locations may be distinguished by their behavior over time in multiple image sets: small ejecta from the Sun are known to follow radial trajectories, and the ghost location follows a clearly nonphysical trajectory (DeForest et al., 2013b; Gibson et al., 2025).
@@
Figure 19 illustrates simple 3D inversion, applicable to compact features that are well isolated and fully background subtracted. Reading off the polarization ratio yields an estimate of the $\tau$ angle (complement of the scattering angle $\chi$). From that and the elongation $\varepsilon$ of the feature, one may infer the out-of-sky-plane angle $\xi$, as shown in Figure 18. This compact-feature case, for features at least several $r_{\odot}$ from the Sun itself, has been developed analytically by several authors including DeForest et al. (2013b). This compact-feature approximation may be used even for portions of large features (whose size is comparable to the distance from the observer or the Sun), because the 2-D leading edge of any large feature must necessarily be spatially compact along the line of sight (Howard et al., 2013). However, all coronal features exist against the background of myriad other K corona (or solar wind) features along each line of sight, and that “K background” must be considered and removed in addition to the backgrounds from stray light, F corona, and stellar backgrounds described in Section 6.4 (e.g., van de Hulst, 1950; DeForest et al., 2017).